### v2

@@ 3.3.3 Spectral Analysis: CTSS 2
The continuum in the nuclear spectra contains absorption signatures (H and He), which we attribute to the central star. We also detect an emission doublet of C IV 5801, 5812 Å, typical of many hot stars. To infer the stellar parameters we performed a spectral fit. We computed a series of non-local thermodynamic equilibrium models of the type introduced by Werner et al. (2019) using the Tübingen Model-Atmosphere Package (Werner et al., 2003). They are plane-parallel and in radiative and hydrostatic equilibrium. They are composed of hydrogen and helium while carbon and nitrogen were included as trace elements keeping fixed the atmospheric structure. We assumed solar composition. Models were computed in the range = 50–90 kK and = 3.9–4.8. The best-fit parameters were determined by eye and found to be kK and (bottom three panels in Figure 10 shows the fit to the normalized spectrum). Using evolutionary tracks by Miller Bertolami (2016) this corresponds to stellar mass, radius, and luminosity of , , and . The corresponding age of the central star is inferred to be years. This makes it a candidate for one of the youngest post-AGB phase CSPN. We comment here on the possible “contamination” from any nebular continuum. We do not expect significant contamination from the faint diffused nebula, but any continuum associated with the central emitting core can potentially dilute the absorption lines, rendering the spectral fit suboptimal. As this emission is mostly unresolved, appropriate modelling of the emission lines is needed to correct for any such contamination. Such a detailed study is beyond the scope of this paper.
The Gaia parallax distance to CTSS 2 is kpc. The analysis of Bailer-Jones et al. (2021) yields a slightly different distance estimate of kpc. We use the inferred stellar parameters to get an independent estimate of the distance to CTSS 2. We first employ a simple manual comparison of the model to the observed spectrum, by varying the reddening and the distance. A good visual fit was obtained with (i.e. ) and kpc (see inset in the top panel of Figure 10). The distance is consistent with the Gaia distances.
@@
From the HST image in Stanghellini et al. (2016), we estimate the angular diameter of the nebula to be around . Assuming a distance to the nebula of kpc, gives the physical radius of the nebula is pc. This gives an average expansion speed of . This is marginally higher than the average expansion velocities of PNe (, Pereyra et al. 2013; Jacob et al. 2013).
@@ 3.3.4 Spectral Analysis: K 3-5
For the nucleus, however, we derive a significantly lower of . For any temperature, this yields , significantly higher than the inferred nebular density, in line with it being a high-density emission core. We note here that weak [S II] lines are also detected in the nuclear spectrum, bearing a similar intensity ratio as the nebula. This is inconsistent with the low value. We thus conclude that this is just a contamination in the nuclear aperture from the overlapping nebula. Unlike CTSS 2, no absorption signature from the central star is detected in the nuclear spectrum preventing any parameter estimation of the CSPN. However, certain emission lines might arise from the central stellar system. For example, we detect C IV 5801, 5812 Å emission line in the nuclear spectrum (absent in the nebular spectrum). This doublet is often associated with either the photosphere of a hot CSPN (like CTSS 2) or an irradiated companion. If the CSPN is an emission line star, H or He emissions from the CSPN can contribute to the observed spectrum. However, further studies are needed to disentangle the nebular and the stellar origin.
@@
3.3.5 Spectral Energy Distribution: Infrared Excess
To look for the presence of dust in these systems, we investigate their spectral energy distributions (SED). The result is presented in Figure 13. Some scatter is evident in the data, especially for K 3-5, which is possibly due to the extended nature of the objects leading to suboptimal photometry. Nevertheless, the presence of significant infrared excess in both the objects is evident.
For a more quantitative estimate, we perform multi-component blackbody fit to the data points with wavelengths . For CTSS 2, a reasonable fit was obtained with a two-component fit. The stellar component was subtracted and distance of kpc was assumed. For K 3-5, we assumed the Gaia eDR3 parallax distance of pc, and a reasonable fit was only observed with three components. The fit parameters of the different components are presented are annotations in the same Figure 13.
@@ 3.3.6 Discussion
The observed photometric evolution can result from several different possible scenarios. Firstly, the observed light curve can be a part of a longer periodic variability due to binarity. Given we do not see the turnover in the light curve over the current baseline (except possibly in WISE data for PN K 3-5), the lower bound to the period is years. The observed variability, thus, would have to be from the irradiation effect. However, such a large amplitude variation is unlikely at such a wide orbit, making this scenario less probable.
Motivated by the inferred stellar parameters of CTSS 2, we consider a rare but very important scenario: late thermal pulses (LTP). LTP designates the occurrence of a He flash in the post-AGB phase, but prior to entering the white dwarf cooling track. The subsequent evolution depends strongly on the conditions at the He flash. But, a typical post-LTP evolution involves expansion of the star (and thus increase in luminosity) along with decrease in temperature as it re-enters the AGB phase (see for example the theoretical tracks discussed in Lawlor 2023). A past example includes FG Sge, which showed extensive photometric brightening (before it turned into an R Coronae Borealis star), at a very similar rate as CTSS 2 (see, for example, Figure 3 in van Genderen & Gautschy 1995). The cooling of the star can also explain the strong reddening observed in the light curves of both CTSS 2 and K 3-5. Note that an LTP model of a 0.625 star was used to explain the evolution of FG Sge (Jeffery & Schönberner, 2006), which agrees very well with the inferred mass of the CSPN of CTSS 2. Additionally, an LTP candidate with parameters very close to that of CTSS 2 was reported in Werner et al. (2025, GSC08265). However, unlike CTSS 2, the spectrum of GSC08265 shows extensive signatures of C and O, indicative of the dredge-up. But, LTP models indicate that the dredge-up occurs only much later (post or near attainment of the minimum temperature, Jeffery & Schönberner 2006). In most cases of LTP (like LTP type III, Lawlor 2023), the photospheric composition remains nearly unchanged during the initial post-LTP phases of the star. CTSS 2 stands as a candidate to an LTP in this stage. K 3-5 might also be a LTP candidate, given the similarity in light curve behavior. But without any spectroscopic identification of the CSPN, the inference is more uncertain.
@@
We now explore other possible scenarios. One object showing a similar long-term photometric evolution is IC 4997. It shows decades-long photometric evolution comprising of both dimming and brightening episodes (Aller & Liller, 1966; Kostyakova & Arkhipova, 2009; Arkhipova et al., 2020). It is a young planetary nebula. The CSPN has been classified to be a possible weak emission-line star (WELS) primarily owing to the presence of C IV 5810, 5812 Å and C III 4650 Å emission lines (note that the former is present in both CTSS 2 and K 3-5). The current understanding attributes the observed photometric behavior to changing wind/mass-loss episodes. As discussed, both objects show evidence of wind through broad H lines. Additionally, significant photometric reddening during brightening has been observed in IC 4997 (Arkhipova et al., 2020), similar to both these objects. Alongside, IC 4997 also shows spectroscopic evolution with changing line ratios indicating changing and over time, also attributed to wind and shocks. This warrants a careful comparison of the current and the archival spectra111111Comparison of the current and past spectra indeed indicates spectroscopic evolution of both CTSS 2 and K 3-5. The spectra of both of them on HASH (and in Stanghellini et al. (2016) for CTSS 2) has H strength much less than the neighboring [O III] lines, and a weak [O III] 4363 Å. This might indicate an increase in associated with the brightening. But we do not establish it with confidence because of a lack of knowledge about the details of the archival spectra (trace selection, nucleus vs nebula).. Because of these reasons, IC 4997 also has a very dense nebula, similar to the core of both these objects.
Long-term dust activity can also lead to photometric evolution. A specific example is M 2-29 (Hajduk et al., 2008; Miszalski et al., 2011b). Both CTSS 2 and K 3-5 bear many similarities to M 2-29: broad H wings indicating wind activity, high-density emission-line core, and presence of emission lines from the CSPN. The presence of extensive dust structures in both objects is clearly evident from the SED. Brightening in such cases results from recovery from heavy dust obscuration. However, we expect blue-ing instead of reddening in such cases. Thus, this scenario appears unlikely though longer photometric monitoring is desired.
@@
The high-density emission core in these nebulae is itself an intriguing aspect. This is the property of the so called EGB 6-type PNe (Liebert et al., 2013) bearing Compact Emission Knots (CEKs), hypothesized to be an irradiated dense accreted envelope around a wide-orbit companion to the CSPN white dwarf. Thus, this may be an indirect indication of the binarity of the CSPNe. For K 3-5, the bipolar nature of the nebula also supports multiplicity of the CSPN. Another explanation involving a companion might be a symbiotic interaction, where such high density cores are common. The lack of the spectral signatures of a Mira giant in the red end (see the DBSP spectrum with a better red coverage, Figure C1), however, speaks against them being the more common D-type symbiotic system. They can be the less common S or D’-type symbiotics. Presence of these systems inside PNe are rare. Other confirmatory signatures of symbiotic behavior, like the Raman-scatter O VI emissions, are also not evident. Further multi-wavelength spectroscopic observations are required to understand these objects.
@@ 3.4 LoTr 1
LoTr 1 is a True PN in HASH and has a diameter of 145′′. It’s CSPN, AB Lep, is a known variable.
@@ 3.4.1 Light Curves
We now note that LoTr 1 is similar to LoTr 5, which consists of an unresolved hot (pre-)white dwarf and rapidly rotating magnetically active G5 giant pair. For the latter system Kővári et al. (2019) uncovered a long-term photometric variability which appears in phase with the orbital radial velocity of the binary ( yr). Thus, one can also speculate that the long-term variability in LoTr 1 might also reflect the orbital period of the binary system. Thus, unlike Martínez et al. (2022), we propose that this is a wide binary with a rapidly rotating giant companion.
3.4.2 HST/COS Spectroscopy and Spectral Energy Distribution
@@
In order to estimate the radii of both stars we carried out a two component fit to the spectral energy distribution (SED) of LoTr 1, employing the SED fitting routine described in Heber et al. (2018) and Irrgang et al. (2021). For the cool star we used PHOENIX models calculated by Husser et al. (2013) and for the hot central star the metal free model grid computed by Reindl et al. (2023). We fixed the of the central star to 105 000 K (i.e. to the value that was derived for UCAC2 46706450 by Werner et al. 2020), and then let the let the angular diameter, reddening, surface ratio of both stars, as well as the of the cool companion vary freely. By that we derive a reddening of mag121212Fitzpatrick et al. (2019) employs , which is the monochromatic equivalent of usual , using the wavelengths 4400 Å and 5500 Å, respectively. For high effective temperatures such as for the stars in our sample is identical to . and a radius of 0.039 for the hot central star (identical to UCAC2 46706450), indicating that it is just about to enter the white dwarf cooling sequence. For the cool giant star we derive K and a radius of 9.7 . These parameters are broadly consistent with a subgiant star of mass in the range of ,131313We compared the stellar parameters to Modules for Experiments in Stellar Astrophysics (MESA, Paxton et al. 2011, 2013, 2015) evolutionary tracks queried from MESA Isochrones and Stellar Tracks (MIST, Dotter 2016; Choi et al. 2016). Owing to the preliminary nature of the analysis we do not show the comparison. This will be presented elsewhere with spectroscopic observations. which is also very similar to the companion of UCAC2 46706450. With the 6.48 day rotational period, this implies a surface velocity of .
3.4.3 Discussion: Growing Evidence of Interaction in Wide Binaries during AGB?
We briefly discuss here a growing class of objects of which LoTr 1 is a likely members: a hot white dwarf with a rapidly rotating giant companion at a wide orbit. In PN literature, such systems are known as Abell 35 type CSPNe (after the first such system, Jacoby 1981; Jasniewicz & Acker 1988). There are several such CSPNe known (see Bond & Zeimann 2024 and their Section 8 for a concise account of such CSPNe). We note here that WeSb 1, the recently discovered CSPN with orbiting dust and debris disk (Bhattacharjee et al., 2025; Budaj et al., 2025) is also likely to be a similar system. Recent high resolution optical spectrum has shown that the subgiant companion of WeSb 1 is rapidly rotating with (details to be presented elsewhere, Bhattacharjee et al. in preparation). A likely formation channel of these systems is mass transfer from the white dwarf progenitor AGB to the companion which spins it up (see, for example, Jeffries & Stevens 1996). These systems, thus, potentially hold valuable information about binary interactions at wide orbits.
The occurrence of such system will be less likely down the white dwarf cooling track as the companion is expected to eventually spin down. The discovery of UCAC2 46706450, however, shows that such systems can be found even after the associated PN has faded away. Thus, we speculate that a population of such sources exist outside of PNe and are yet to be found and studied.
@@ 3.5 The other long-timescale variables
We first discuss the True PNe. PN K 3-20 shows a brightening event, with a larger amplitude in the ZTF band. Tan 2 shows a singular eclipse-like dip in the band. Unfortunately, the band data for this object did not pass the photometric quality cuts. The light curve of WHTZ 3, albeit noisier, shows low amplitude long-term modulation. The possible causes for variability of these objects are uncertain. The observed light curves can be a part of even longer timescale periodic variability associated with binarity. For example, the behavior in PN K 3-20 may result from irradiation effect in a long-period wide binary (similar to NGC 6833). Further long-term photometric monitoring is needed to test these scenarios. Concerning the remaining two objects, Kn J1857 shows a gradual dimming, by in flux, with similar amplitude in both the ZTF bands. The variability in IPHAS J1927, on the other hand, is more irregular. Similar kinds of photometric behaviors are often seen in symbiotic systems or cataclysmic variables (CVs). Being only a ‘Possible PN’, spectroscopic data is necessary to determine the true nature of this object.
@@ 4 Summary and Conclusions
This is the second of a series of two survey papers undertaking a systematic study of the variability of the central stars of planetary nebulae (CSPNe) using the optical light curves from the Zwicky Transient Facility (ZTF). In Paper I, we applied appropriate variability metrics to identify significantly variable sources. This paper focuses on 11 objects among them, which show long-timescale variability. For six of these objects, we present follow-up spectroscopic/photometric observations and provide extensive discussion, which is expected to serve as the first step in understanding the objects. We also present the remaining six objects and encourage follow-up studies by the community.
@@
The first object we discuss is NGC 6833 (Section 3.1). It shows ‘triangle-shaped’ brightening in the ZTF , , and WISE W1 and W2 bands, occurring at a period of days. The more intriguing aspect is the anticorrelated behavior in the ZTF -band, which shows shallow dips nearly at the same epochs as the peaks in the redder bands. We speculate that this might be a long-period eccentric binary, with the photometric period corresponding to the orbital period. This makes it the candidate for the fourth to the longest orbital period inside a PN. Though the irradiation effect can possibly explain the brightenings, the simultaneous dips in -band is puzzling. We speculate this might be due to dust production through certain stellar interactions at orbital periastron.
@@
Long-periodic modulation was also detected in the light curves of NGC 6905 and Kn 26 (Section 3.2. The former showed a significant peak at days and the latter at days. Secondary significant peaks were also detected at days and days were also detected. The modulation appears broadly sinusoidal. This makes both the objects candidates for long-period binary inside CSPNe. Their periods (as well as that of NGC 6833) are short enough for them to be post-CE systems, making them candidates for post-CE wide binaries. Kn 26 displays an additional periodicity at 1.18 hr, which we verify through follow-up photometry. The origin of this period is uncertain. It can be of close binary (in which case the interpretation of the long period as a wide orbit binary makes it a candidate triple system) or pulsational origin.
@@
We present CTSS 2 and PN K 3-5, which show slow and long-term brightening (Section 3.3). The objects brightened by almost a factor of in the ZTF -band and WISE bands, but much less so (factor of ) in the -band, thus undergoing significant reddening. Spectroscopy of both the nucleus and the nebula shows that both objects host high-density () emission-line cores. Thus they are EGB 6-type PNe where accreted ejecta on a wide companion is irradiated, or symbiotic systems. The optical spectrum of CTSS 2 displays stellar signatures. A spectral fit yields a temperature of kK and a stellar mass of . This translates to a post-AGB age of years, making it one of the youngest CSPN known. Unfortunately, the spectrum of K 3-5 does not show adequate stellar features preventing any spectroscopic analysis of the pre-white dwarf. We speculate that these might be late thermal pulse events. In this case, the light curves suggest that the stars are cooling down and expanding, moving back to another AGB phase. This is a very rare occurrence and would indicate the success of long-baseline surveys like ZTF to discover these objects. Other possible explanations include wind/dust activity of the CSPN system.
@@
We present new results on LoTr 1 (Section 3.4), which is a known variable showing a long-period modulation at years and a shorter period at days. The latter is attributed to the rotational modulation of a fast-rotating cool giant companion. The longer period might be the orbital period. Ultraviolet spectroscopy with HST shows that the hot component in the system is a spectroscopic twin of the extremely hot white dwarf in UCAC2 46706450, with a temperature of kK. This hints at the presence of a whole class of objects with fast-rotating giant companions to very hot white dwarfs. Further studies are underway and will be presented elsewhere. A fit to the SED reveals the stellar parameters of the giant to be K and radius of .
@@
Finally, we briefly discuss the observed variability in the remaining five objects (Section 3.5). They display a variety of photometric behaviors: long timescale dimming, single brightening event, single dip event, or more erratic variability. Follow-up studies are necessary to understand these systems.
@@
Overall, this study shows the effectiveness of long-baseline surveys in uncovering the variety of photometric behaviors observed CSPNe. Efforts need to be made to not only understand the objects individually through follow-up observations but, if possible, to establish evolution links to different classes of objects. For example, the peculiar photometric behavior of NGC 6833 may shed significant light on the formation and evolution of post-CE wide binary systems. Note here that ZTF can only access the north sky. The majority of the PNe are in the southern sky, near the galactic center. Upcoming surveys like LSST will observe these targets, and we expect to find several more such intriguing systems. Simple variability metrics and metric spaces (as demonstrated in Appendix A) will significantly boost such discoveries.

### v3

@@ 3.3.3 Spectral Analysis: CTSS 2
The continuum in the nuclear spectra contains absorption signatures (H and He), which we attribute to the central star. We also detect an emission doublet of C IV 5801, 5812 Å, typical of many hot stars. To infer the stellar parameters we performed a spectral fit. We computed a series of non-local thermodynamic equilibrium models of the type introduced by Werner et al. (2019) using the Tübingen Model-Atmosphere Package (Werner et al., 2003). They are plane-parallel and in radiative and hydrostatic equilibrium. They are composed of hydrogen and helium while carbon and nitrogen were included as trace elements keeping fixed the atmospheric structure. We assumed solar composition. Models were computed in the range = 50–90 kK and = 3.9–4.8. The best-fit parameters were determined by eye and found to be kK and (bottom three panels in Figure 7 shows the fit to the normalized spectrum). Using evolutionary tracks by Miller Bertolami (2016) this corresponds to stellar mass, radius, and luminosity of , , and . The corresponding age of the central star is inferred to be years. This makes it a candidate for one of the youngest post-AGB phase CSPN. We comment here on the possible “contamination” from any nebular continuum. We do not expect significant contamination from the faint diffused nebula, but any continuum associated with the central emitting core can potentially dilute the absorption lines, rendering the spectral fit suboptimal. As this emission is mostly unresolved, appropriate modeling of the emission lines is needed to correct for any such contamination. Such a detailed study is beyond the scope of this paper.
The Gaia parallax distance to CTSS 2 is kpc. The analysis of Bailer-Jones et al. (2021) yields a slightly different distance estimate of kpc. We use the inferred stellar parameters to get an independent estimate of the distance to CTSS 2. We first employ a simple manual comparison of the model to the observed spectrum, by varying the reddening and the distance. A good visual fit was obtained with (i.e. ) and kpc (see inset in the top panel of Figure 7).101010All dust extinction corrections in this work have been performed using the dust_extinction (Gordon, 2024a, b) implementation of the Gordon et al. (2023) galactic extinction models (Decleir et al., 2022; Gordon et al., 2021; Fitzpatrick et al., 2019a; Gordon et al., 2009). The distance is consistent with the Gaia distances.
@@
From the HST image in Stanghellini et al. (2016), we estimate the angular diameter of the nebula to be around . Assuming a distance to the nebula of kpc, gives the physical radius of the nebula is pc. This gives an average expansion speed of . This is marginally higher than the average expansion velocities of PNe (, Pereyra et al. 2013; Jacob et al. 2013), but consistent within the error limits.
@@ 3.3.4 Spectral Analysis: K 3-5
For the nucleus, however, we derive a significantly lower of . For any temperature, this yields , significantly higher than the inferred nebular density, in line with it being a high-density emission core. We note here that weak [S II] lines are also detected in the nuclear spectrum, bearing a similar intensity ratio as the nebula. This is inconsistent with the low value. We thus conclude that this is just a contamination in the nuclear aperture from the overlapping nebula. Unlike CTSS 2, no absorption signature from the central star is detected in the nuclear spectrum preventing any parameter estimation of the CSPN. However, certain emission lines might arise from the central stellar system. For example, we detect C IV 5801, 5812 Å emission line in the nuclear spectrum (absent in the nebular spectrum). This doublet is often associated with either the photosphere of a hot CSPN (like CTSS 2) or an irradiated companion.
@@ 3.3.5 Discussion
CTSS 2 is a good candidate for a late thermal pulse (LTP) scenario, in which the star experiences a late He-shell flash shortly after leaving the AGB (e.g. Lawlor 2023). LTPs are expected to occur frequently, and lead to an accelerated evolution of the star. Thus far, only two such stars have been ”caught in the act”: SAO 244567 was observed during the initial rapid heating and contraction, which was followed around 2002 by a rapid cooling and expansion of the star Reindl et al. (2014, 2017). The other example is FG Sge which was observed only during the cooling an expansion phase, returning back to the AGB, where it turned into a H-deficient now an RCB star (Jeffery & Schönberner, 2006). We stress, that the brightening rate of CTSS 2 is very similar to that of FG Sge (see, for example, Figure 3 in van Genderen & Gautschy 1995), and the strong reddening observed in the light curves of CTSS 2 suggests that the star is currently cooling. (Jeffery & Schönberner, 2006) explained the evolution of FG Sge with a 0.625 LTP model, which agrees very well with our inferred mass of the CSPN of CTSS 2. We also note that the H-rich surface composition of CTSS 2 is what is expected from LTP models before the star manages to evolve back to the AGB (e.g. Lawlor 2023) and what was observed for SAO 244567 and FG Sge. K 3-5 might also be a LTP candidate, given the similarity in light curve behavior. But without any spectroscopic identification of the CSPN, the inference is more uncertain.
We now consider a few other possible scenarios. Firstly, we deem it unlikely that the observed light curves are part of a periodic binary signal. The lack of turnover in the light curves implies years. At such a wide orbit, neither irradiation nor ellipsoidal variability can be strong enough to cause the observed brightening. The former might be significant for a highly eccentric orbit, but the duration of the periastron (thus the irradiation effect) then would be much smaller. Thus, we deem this scenario highly unlikely.
@@
The two systems can be similar to IC 4997 which shows decades-long photometric evolution comprising of both dimming and brightening episodes (Aller & Liller, 1966; Kostyakova & Arkhipova, 2009; Arkhipova et al., 2020). The cause for the variability is yet unclear. The other possibility is long term dust activity, as witnessed in M 2-29 (Hajduk et al., 2008; Miszalski et al., 2011b). The spectral energy distributions of both CTSS 2 and K 3-5 indeed indicate presence of a large amount of dust (see Appendix F). Long term photometric and spectroscopic111111Comparison of the current and past spectra indeed indicates spectroscopic evolution of both CTSS 2 and K 3-5. The spectra of both of them on HASH (and in Stanghellini et al. (2016) for CTSS 2) has H strength much less than the neighboring [O III] lines, and a weak [O III] 4363 Å. This might indicate an increase in associated with the brightening. But we do not establish it with confidence because of a lack of knowledge about the details of the archival spectra (trace selection, nucleus vs nebula). monitoring is needed for a more concrete understanding of these systems.
@@
The high-density emission cores in both the systems warrant a separate mention. This is a property of the so-called EGB 6-type PNe (Liebert et al., 2013) bearing Compact Emission Knots (CEKs), hypothesized to be an irradiated dense accreted envelope around a wide-orbit companion to the CSPN white dwarf. Thus, this may be an indirect indication of the binarity of the CSPNe. For K 3-5, the bipolar nature of the nebula also supports multiplicity of the CSPN. Another explanation involving a companion might be a symbiotic interaction, where such high density cores are common. The lack of the spectral signatures of a Mira giant in the red end (see the DBSP spectrum with a better red coverage, Figure E1), however, speaks against them being the more common D-type symbiotic system. They can be the less common S or D’-type symbiotics. Presence of these systems inside PNe are rare. Confirmatory signatures of symbiotic behavior, like the Raman-scattered O VI emissions, are not seen in the spectra. But this does not rule out symbiotic interaction as such signatures are highly variable and not present in many systems.
@@ 3.4 LoTr 1
LoTr 1 is a True PN in HASH and has a diameter of 145′′. Its CSPN, AB Lep, is a known variable.
@@ 3.4.1 Light Curves
We note that LoTr 1 is similar to LoTr 5, which consists of an unresolved hot (pre-)white dwarf and rapidly rotating magnetically active G5 giant pair. For the latter system Kővári et al. (2019) uncovered a long-term photometric variability which appears in phase with the orbital radial velocity of the binary ( yr). Thus, one can likewise speculate that the long-term variability in LoTr 1 might reflect the orbital period of the binary system. Unlike Martínez et al. (2022), we propose that this is a wide binary with a rapidly rotating giant companion.
3.4.2 HST/COS Spectroscopy and Spectral-Energy Distribution
@@
In order to estimate the radii of both stars we carried out a two component fit to the spectral energy distribution (SED) of LoTr 1, employing the SED fitting routine described in Heber et al. (2018) and Irrgang et al. (2021). For the cool star we used PHOENIX models calculated by Husser et al. (2013) and for the hot central star the metal free model grid computed by Reindl et al. (2023). We fixed the of the central star to 105 000 K (i.e. to the value that was derived for UCAC2 46706450 by Werner et al. 2020), and then let the let the angular diameter, reddening, surface ratio of both stars, as well as the of the cool companion vary freely. By that we derive a reddening of mag121212Fitzpatrick et al. (2019b) employs , which is the monochromatic equivalent of usual , using the wavelengths 4400 Å and 5500 Å, respectively. For high effective temperatures such as for the stars in our sample is identical to . and a radius of 0.44 for the hot central star (identical to UCAC2 46706450), indicating that it is just about to enter the white dwarf cooling sequence. For the cool giant star we derive K and a radius of 9.7 . These parameters are broadly consistent with a subgiant star of mass in the range of ,131313We compared the stellar parameters to Modules for Experiments in Stellar Astrophysics (MESA, Paxton et al. 2011, 2013, 2015) evolutionary tracks queried from MESA Isochrones and Stellar Tracks (MIST, Dotter 2016; Choi et al. 2016). Owing to the preliminary nature of the analysis we do not show the comparison. This will be presented elsewhere with spectroscopic observations. which is also very similar to the companion of UCAC2 46706450. With the 6.48 day rotational period, this implies a surface velocity of .
@@ 3.5 The other long-timescale variables
We first discuss the True PNe. K 3-20 shows a brightening event, with a larger amplitude in the ZTF band. Tan 2 shows a singular eclipse-like dip in the band. Unfortunately, the band data for this object did not pass the photometric quality cuts. The light curve of WHTZ 3, albeit noisier, shows low amplitude long-term modulation. The possible causes for variability of these objects are uncertain. The observed light curves can be a part of even longer timescale periodic variability associated with binarity. For example, the behavior in K 3-20 may result from irradiation effect in a long-period wide binary (similar to NGC 6833). Further long-term photometric monitoring is needed to test these scenarios. Concerning the remaining two objects, Kn J1857 shows a gradual dimming, by in flux, with similar amplitude in both the ZTF bands. The variability in IPHAS J1927, on the other hand, is more irregular. Similar kinds of photometric behaviors are often seen in symbiotic systems or cataclysmic variables (CVs). Being only a “Possible PN,” spectroscopic data are necessary to determine the true nature of this object.
@@ 4 Summary and Conclusions
This is the second in a series of papers undertaking a systematic study of the variability of central stars of planetary nebulae (CSPNe), using optical light curves from the Zwicky Transient Facility (ZTF). In Paper I, we applied appropriate variability metrics to identify significantly variable sources. The present paper reports the 11 objects which show long-timescale variability. For six of these objects, we present follow-up spectroscopic/photometric observations and provide extensive discussion, which is expected to serve as the first step in understanding the objects. We also present the remaining six objects and encourage follow-up studies by the community.
@@
The first object we discuss is NGC 6833 (Section 3.1). It shows ‘triangle-shaped’ brightening in the ZTF , , and WISE W1 and W2 bands, occurring at a period of days. The more intriguing aspect is the anti-correlated behavior in the ZTF -band, which shows shallow dips nearly at the same epochs as the peaks in the redder bands. We think that this is a long-period eccentric binary, with the photometric period corresponding to the orbital period. This makes it the candidate for the fourth-longest binary orbital period inside a PN. Though the irradiation effect can possibly explain the brightening, the simultaneous dips in -band is puzzling. Dust occultation and extinction can potentially result in such a feature.
@@
Long-periodic modulation was also detected in the light curves of NGC 6905 and Kn 26 (Section 3.2). The former showed a significant peak at days and the latter at days. Secondary significant peaks were also detected at days and days. The modulation appears broadly sinusoidal. This makes both the objects candidates for long-period binary CSPNe. Kn 26 displays an additional periodicity at 1.18 hr, which we verify through follow-up photometry. The origin of this period is uncertain. It can be due to a close binary (unlikely) or rotational origin.
@@
We present CTSS 2 and K 3-5, which show slow and long-term brightening (Section 3.3). The objects brightened by almost a factor of in the ZTF -band and WISE bands, but much less so (factor of ) in the -band, thus undergoing significant reddening. Spectroscopy of both the nucleus and the nebula shows that both objects host high-density () emission-line cores. Thus they are either EGB 6-type PNe with compact emission knots, or symbiotic systems. The optical spectrum of CTSS 2 displays stellar absorption signatures. A spectral fit yields a temperature of kK and a stellar mass of . This translates to a post-AGB age of years, making it one of the youngest CSPNe known. Unfortunately, the spectrum of K 3-5 does not show adequate stellar features, thus preventing any spectroscopic analysis of the (pre-)white dwarf. We think that these might be late thermal pulse events. In this case, the light curves suggest that the stars are cooling down and expanding, moving back to another AGB phase. This is a very rare occurrence and would indicate the success of long-baseline surveys like ZTF to discover these objects. Other possible explanations include wind/dust activity of the CSPN system.
@@
We present new results on LoTr 1 (Section 3.4), which is a known variable showing a long-period modulation at years and a shorter period at days. The latter is attributed to the rotational modulation of a fast-rotating cool giant companion. The longer period might be the orbital period. Ultraviolet spectroscopy with HST shows that the hot component in the system is a spectroscopic twin of the extremely hot white dwarf in UCAC2 46706450, with a temperature of kK. A fit to the SED reveals the stellar parameters of the giant to be K and radius of .
@@
Finally, we briefly discuss the observed variability in the remaining five objects (Section 3.5). They display a variety of photometric behaviors: long-timescale dimming, single brightening event, single dip event, or more erratic variability. Follow-up studies are necessary to understand these systems.
@@
Overall, this study shows the effectiveness of long-baseline surveys in uncovering the variety of photometric behaviors observed in CSPNe. Efforts need to be made not only to understand the objects individually through follow-up observations, but, if possible, to establish evolutionary links to different classes of objects. For example, the peculiar photometric behavior of NGC 6833 may shed significant light on the formation and evolution of post-CE wide binary systems. Note here that ZTF can only access the northern sky. The majority of PNe are in the southern sky, near the Galactic center. Upcoming surveys like LSST will observe these targets, and we expect to find several more such intriguing systems. Simple variability metrics and metric spaces (as demonstrated in Appendix A) will significantly boost such discoveries.